OpenStudy (scienceteen8):

i will give a medal A number is written with the following factorization: 2^3 × 4 × 5. Is this factorization a prime factorization? Explain why or why not. If it is not correct, give the correct prime factorization of the number.

OpenStudy (alexandervonhumboldt2):

4=2*2. if even 1 element of the factorization has dividers then it is not a prime factorization. 2^3*2*2*5=2^5*5 it should be

OpenStudy (scienceteen8):

Thnxs!! :)